OpenStudy (anonymous):

q = 13/2 + 47/6 then find common denominator

OpenStudy (texaschic101):

common denominator for 6 and 2 is 6. q = 13/2 + 47/6 q = 39/6 + 47/6...now just add I believe you can finish this problem now....

OpenStudy (aravindg):

@elizabethnal94 gt it?

OpenStudy (anonymous):

i think so

OpenStudy (texaschic101):

what answer did you get ?

OpenStudy (aravindg):

:) \[\huge \color{green}{\text{WELCOME TO OPENSTUDY!}}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

the quiestion is where did you get the 39?

OpenStudy (texaschic101):

13/2 +47/6 (the common denominator is 6) To do this problem, you have to make both denominators 6. The 47/6 is ok, but you have to make the 13/2 have a denominator of 6. 13/2....how many times will 2 go into 6 ? It will go into 6, three times. Now you take your 3 and multiply it by the numerator, which is 13, and 13 x 3 = 39.Now you put your 39 over the new denominator. Your problem is now 39/6 + 47/6. The denominators have to be the same in order to add them. Do you understand ?